OpenStudy (anonymous):

A plane left its base at noon and flew 800 miles directly against the stream with an air speed of 750mph(speed of plane in still air). It then returned directly to its base arriving at 2:24pm. What was the speed of the jet stream(speed of the air)?

OpenStudy (anonymous):

options?

OpenStudy (anonymous):

(750-x)*t1 =800 (750+x)*t2 = 800 t1+t2 = 2.4

OpenStudy (anonymous):

there are no options